# Beamforming and Waveform Optimization for RF Wireless Power Transfer with Beyond Diagonal Reconfigurable Intelligent Surfaces

### v1

@@ III-B SDR-based Beamforming for Fully Connected BD-RIS
Algorithm 2 describes the proposed SDR-based method for beamforming and waveform optimization for fully connected BD-RIS. First, the optimization variables are initialized. Then, the waveform and scattering matrix are optimized in an alternative fashion through lines 7-17. Specifically, beamforming is done by solving (25) iteratively in lines 9-14, followed by iterative waveform optimization using Algorithm 1. Finally, Algorithm 3 is utilized to construct a feasible solution $\mathbf{\Theta}^{*}$ based on the characteristics of the obtained solution $\mathbf{\Theta}$. First, the SVD of $\mathbf{\Theta}$ is computed to obtain a unitary matrix $\mathbf{Q}$. Then, random phase shifts are generated for $K$ iterations to construct new $\boldsymbol{\Sigma}^{\prime}$ matrices and their corresponding feasible solution $\mathbf{\Theta}$. Finally, the constructed $\mathbf{\Theta}$ with the best $i_{dc}$ is selected.
III-C Iterative Beamforming Optimization
@@
  • Proof.
    The proof is provided in Appendix C.
By leveraging Proposition 2, one can linearize $h_{n}$ using the Neumann approximation and iteratively improve the approximation. Thus, by defining an auxiliary variable $\mathbf{\Omega}$ as the increment to the impedance matrix $\mathbf{Z}^{(l)}$ at the $l$th iteration of the Neumann approximation, we can write
@@
By leveraging the same matrix properties provided in Appendix B, we proceed by rewriting
$h_{n}$ $\approx{\mathbf{a}_{n}^{(l)}}^{T}(\mathbf{I}_{N}-\mathbf{\Omega}\mathbf{A}^{(l)})\mathbf{b}_{n}^{(l)}$
@@
Proposition 3 and Lemma 2 ensure that substituting (30) into (26a) leads to maximizing a convex objective function w.r.t. $\boldsymbol{\omega}$, which can be approximated similar to the waveform optimization procedure for fixed $h_{n}$. Specifically, the convex objective can be approximated using its first-order Taylor’s expansion at each local point and iteratively optimized. Thus, we proceed by writing the first-order Taylor’s coefficient as
@@
where (34c) ensures that $\mathbf{\Omega}$ has the same zero entries as $\mathbf{Z}$ depending on the structure of BD-RIS. Note that for a fully-connected BD-RIS, constraint (34c) can be removed.
Theorem 2.
The local solution of (34) can be updated using
@@
where $\rho_{\omega}\leq 1$ is a parameter utilized to control the convergence of the solution and
$[\boldsymbol{\omega}^{*}]_{r}$ =$=\begin{cases}0,&r\in\mathcal{D}_{\mathcal{G}},\\ {[\mathbf{u}^{(l)}]_{r}}/{2\nu_{r}^{(l)}},&\text{otherwise},\end{cases}$ (36)
$\nu_{r}^{(l)}$ =$=\begin{cases}0,&r\in\mathcal{D}_{\mathcal{G}},\\ \left\lVert[\mathbf{u}^{(l)}]_{r}\right\rVert/2{\tau^{(l)}},&\text{otherwise}.\end{cases}$ (37)
  • Proof.
    The proof is provided in Appendix D.
We leverage Theorem 2 to solve (34) iteratively. Moreover, it was verified that $\mathrm{Vec}(\mathbf{\Omega}^{(l+1)})=\mathbf{P}\boldsymbol{\omega}^{(l+1)}$; thus, $\mathbf{\Omega}^{(l+1)}=\mathrm{Vec}^{-1}_{M,M}(\mathrm{Vec}(\mathbf{\Omega}^{(l+1)}))$ can be used for updating $\mathbf{Z}$ as
@@
Algorithm 4 illustrates the proposed iterative optimization method relying on alternating optimization, SCA, KKT conditions, and the Neumann approximation for BD-RIS-aided WPT (IT-BDRIS) given $\mathcal{D}_{\mathcal{G}}$. First, we initialize the variables such that the impedance matrix is set to reference impedance value, while (13) is used to initialize $s_{n}$. Then, the alternating optimization is done through lines 6-16. Specifically, each alternating optimization iteration consists of iteratively updating $\mathbf{Z}$ given fixed $s_{n},\forall n$ in lines 8-13, followed by obtaining $s_{n},\forall n$ using Algorithm 1. This procedure is repeated iteratively until convergence.
@@
III-D Complexity Analysis
Herein, we discuss the time complexity of the proposed algorithms in details:
IT-WF: The time complexity at each optimization iteration is $\mathcal{O}(N)$, which corresponds to the computation of $s_{n}$.
SDR-BDRIS: The number of variables in (25) is ${\bar{M}(\bar{M}+1)}/2$ with $\bar{M}=M(M+1)/2$, while the rest of the entries in $\mathbf{X}$ are determined according to the Hermitian structure, and the size of these Hermitian matrix sub-space is ${\bar{M}}^{2}$. Additionally, the number of constraints scales with ${M}$ in (25). For a given accuracy, the complexity of SDP grows at most with $\mathcal{O}(n^{1/2})$, where $n$ is the problem size, determined by the number of constraints and variables [45]. Thus, the SDR-BDRIS converges in polynomial time and its complexity increases with $K$ and $\bar{M}^{2}$ combined with complexity for IT-WF. Thus, it is evident that the complexity of SDR-BDRIS is significantly higher than IT-BDRIS since $\bar{M}^{2}$ can become difficult to handle for large BD-RIS size.
IT-BDRIS: In the most time-consuming case, i.e., fully connected impedance network, $\boldsymbol{\omega}$ includes $\bar{M}$ non-zero entries, leading to $\mathcal{O}(\bar{M})$ complexity for obtaining the solution and the dual variables. Combining this with the negligible complexity for obtaining $\mathcal{Z}^{(l+1)}$ leads to $\mathcal{O}(\bar{M})$ time complexity for each optimization iteration. Combining this with the complexity of IT-WF evinces that the proposed IT-BDRIS is much more efficient than the traditional solver-based SCA.
@@ IV Numerical Analysis
Figure 2: The convergence performance of (a) IT-WF (left), (b) iterative beamforming (middle-left), (c) IT-BDRIS for joint optimization (middle-right), (d) SDR-BDRIS for joint optimization (right). We adopt a random channel realization and assume a fully-connected BD-RIS and several $M$ and $N$ combinations.
In this section, we evaluate the system’s performance in a WiFi-like scenario with $f_{c}=2.4$ GHz, a bandwidth of $\textrm{BW}=10$ MHz, and $\Delta f=\textrm{BW}/N$. We consider the path loss due to large-scale fading at a distance $d$ to be $L_{0}d^{-\kappa}$, where $L_{0}=40$ dB is the reference path loss at 1 m distance and $\kappa=2$ is the exponent. Moreover, the incident and reflected paths are both considered to be 2 m. Finally, the rectifier parameters are selected based on the circuit simulations in [8, 22] for small-signal regime such that $K_{2}=0.17$ and $K_{4}=957.25$.
@@
where $\kappa$ is the Rician factor and the $\mathbf{h}_{n}^{\text{LoS}}$ is modeled using the far-field LoS channel for uniform planar arrays. The NLoS part is modeled with Rayleigh fading considering $L=18$ delay taps with realizations following a circularly symmetric complex Gaussian distribution with a random power $p_{l}$, such that $\sum_{l=1}^{L}p_{l}=1$. Furthermore, we introduce an additional parameter $\alpha$ for the NLoS part such that the coherence bandwidth is $B_{c}=\alpha\textrm{BW}$. Hereby, the delay spread is $\varsigma=1/B_{c}$, while the maximum delay spread is considered to be $\varsigma_{m}=2\varsigma$. Then, the tap delays $t_{l},\forall l$ are generated uniformly spaced up to $\varsigma_{m}$, and the frequency response of each tap at the $n$th sub-carrier is multiplied by its corresponding delay term $e^{j2\pi f_{n}t_{l}}$. Hereby, selecting a small $\alpha$ such that $B_{c}$ is much smaller than BW leads to more frequency selectivity of the channel. Nevertheless, a large $\alpha$ leads to $B_{c}\gg\textrm{BW}$, which means all delay taps can be received within the coherence bandwidth, leading to a frequency-flat channel. The results are averaged over 200 random channel realizations unless otherwise stated.

### v2

@@ III-B SDR-based Beamforming for Fully Connected BD-RIS
Algorithm 2 describes the proposed SDR-based method for beamforming and waveform optimization for fully connected BD-RIS. First, the optimization variables are initialized. Then, the waveform and scattering matrix are optimized in an alternative fashion through lines 7-17. Specifically, beamforming is done by solving (27) iteratively in lines 9-14, followed by iterative waveform optimization using Algorithm 1. Finally, Algorithm 3 is utilized to construct a feasible solution $\mathbf{\Theta}^{*}$ based on the characteristics of the obtained solution $\mathbf{\Theta}$. First, the SVD of $\mathbf{\Theta}$ is computed to obtain a unitary matrix $\mathbf{Q}$. Then, random phase shifts are generated for $K$ iterations to construct new $\boldsymbol{\Sigma}^{\prime}$ matrices and their corresponding feasible solution $\mathbf{\Theta}$. Finally, the constructed $\mathbf{\Theta}$ with the best $i_{dc}$ is selected. Note that the SDP-BDRIS algorithm follows the same procedure as Algorithm 2, but solves (25) instead of (27).
The problems in (25) and (27) are bounded due to the constraints in (21b), (21c), and (25c), along with the rank surrogate constraint (25b) in the SDP-based formulation. The proposed SCA-based beamforming algorithms, based on SDP and SDR, iteratively solve convex approximations of the original non-convex problem (21). At each iteration, both the objective and the non-convex surrogate rank constraint (in the SDP-based case) are approximated using first-order Taylor expansions that satisfy the standard SCA assumptions [scutari2016paralleldistributedmethodsnonconvex, DC_constraint_SCA]. As in Theorem 1, the SDP-based beamforming converges to a stationary point of the original beamforming problem (21). For the SDR-based case, convergence to a stationary point of (21) holds only when the obtained solution of the relaxed problem is rank-1. Therefore, since Algorithm 1 converges to a stationary point of the waveform subproblem (6), the full SDP-BDRIS algorithm yields a stationary point of the joint problem (5). On the other hand, SDR-BDRIS converges to a stationary point of (5) only when the solution to (27) is rank-1. Otherwise, if the relaxed solution is of a higher rank, the rank-1 solution obtained via Algorithm 3 is a feasible approximate solution, and the resulting point is not guaranteed to satisfy the stationary conditions of the original problem (5).
III-C SCA-based Beamforming for Fully Connected BD-RIS
Here, we provide an optimization framework relying only on SCA by directly approximating (16). Specifically, we directly deal with the matrix $\boldsymbol{\Theta}$ without lifting the variable size. Let us proceed by writing the first-order Taylor approximation of $i_{dc}$ w.r.t. $\boldsymbol{\Theta}$ as
$\tilde{i}_{dc}(\boldsymbol{\Theta},\boldsymbol{\Theta}^{(l)})=i_{dc}(\boldsymbol{\Theta}^{(l)})\\ +2\,\Re\!\left\{\operatorname{Tr}\!\left(\big(\nabla_{\boldsymbol{\Theta}}i_{dc}(\boldsymbol{\Theta}^{(l)})\big)^{\!H}\,(\boldsymbol{\Theta}-\boldsymbol{\Theta}^{(l)})\right)\right\}.$ (28)
Moreover, the unitary constraint can be divided into two separate constraints formulated as $\boldsymbol{\Theta}^{H}\boldsymbol{\Theta}\preceq\mathbf{I}_{M}$ and $\boldsymbol{\Theta}^{H}\boldsymbol{\Theta}\succeq\mathbf{I}_{M}$. By applying the Schur complement [boyd2004convex], the former convex constraint can be equivalently expressed as
$\begin{bmatrix}\mathbf{I}_{M}&\boldsymbol{\Theta}^{H}\\ \boldsymbol{\Theta}&\mathbf{I}_{M}\end{bmatrix}\succeq 0,$ (29)
while the latter constraint is nonconvex. To enable tractable optimization, we approximate it via the first-order Taylor expansion around the local point $\boldsymbol{\Theta}^{(l)}$, which leads to the affine surrogate written as [MM_survey]
$g({\boldsymbol{\Theta}^{(l)}},{\boldsymbol{\Theta}})={\boldsymbol{\Theta}^{(l)}}^{H}\boldsymbol{\Theta}+\boldsymbol{\Theta}^{H}\boldsymbol{\Theta}^{(l)}-{\boldsymbol{\Theta}^{(l)}}^{H}\boldsymbol{\Theta}^{(l)}.$ (30)
Since $(\boldsymbol{\Theta}-\boldsymbol{\Theta}^{(l)})^{H}(\boldsymbol{\Theta}-\boldsymbol{\Theta}^{(l)})\succeq 0$, this affine constraint provides a global inner approximation of the original constraint. Hence, the two constraints jointly enforce an approximate unitary constraint within the SCA framework. Note that enforcing these two constraints on the problem may lead to an empty feasible space, thus solver failure. To address this, we introduce a positive semidefinite auxiliary variable $\mathbf{S}$ in the affine surrogate constraint, while penalizing its trace in the objective function. This avoids solver failures by relaxing the unitary constraint to some extent, while the penalty pushes the auxiliary term towards zero, thereby encouraging meeting the constraint throughout the SCA iterations. Moreover, to avoid divergence when the local point $\boldsymbol{\Theta}^{(l)}$ is far from the unitary set, we introduce a trust region that restricts the update $\boldsymbol{\Theta}-\boldsymbol{\Theta}^{(l)}$ within a bounded neighborhood. These two strategies are standard in SCA to guarantee convergence [scutari2016paralleldistributedmethodsnonconvex, boyd2004convex]. Hereby, the local convex problem in the neighborhood of $\boldsymbol{\Theta}^{(l)}$ can be written as
$\operatorname*{maximize}_{\boldsymbol{\Theta},\mathbf{S}\succeq 0}\quad$ $\Omega_{\textrm{sca}}=\tilde{i}_{dc}(\boldsymbol{\Theta},\boldsymbol{\Theta}^{(l)})-\sigma\mathrm{Tr}(\mathbf{S})$ (31a)
subject to $g({\boldsymbol{\Theta}^{(l)}},{\boldsymbol{\Theta}})+\mathbf{S}\succeq\mathbf{I}_{M},$ (31b)
$\left\lVert\boldsymbol{\Theta}-\boldsymbol{\Theta}^{(l)}\right\rVert_{F}\leq\iota,$ (31c)
[…]
Algorithm 4 describes the proposed SCA-based method for beamforming and waveform optimization for fully connected BD-RIS. The overall procedure is similar to Algorithm 2, while beamforming is optimized by iteratively solving (31). Moreover, we use an adaptive penalty coefficient $\sigma$, which is initialized with a small value to preserve feasibility in the early iterations and then gradually increased to enforce the approximate unitary constraint more strictly as the algorithm converges. Furthermore, to avoid infeasible solutions caused by relaxations, we map the final solution using Algorithm 3. The proposed SCA-BDRIS algorithm includes a non-decreasing sequence of the objective values since, in each iteration, the convex surrogate is solved optimally and is tight at the local point. Moreover, every obtained local solution satisfies the KKT conditions of the surrogate problem under standard assumptions [scutari2016paralleldistributedmethodsnonconvex]. In case the obtained solution before mapping is unitary, the final solution is a stationary point of the original problem; otherwise, it is an approximate feasible solution after mapping.
Algorithm 4 SCA-based beamforming and waveform optimization for fully connected BD-RIS (SCA-BDRIS).
1:Input: $\upsilon$, $\sigma$, $\iota$, $\mathbf{h}_{R,n},\mathbf{h}_{I,n},\forall n$Output: $s_{n}^{(l)},\forall n$, $\mathbf{\Theta}^{*}$
2:Initialize: as in Algorithm 2
3:repeat
4:   $\Omega_{\textrm{sca}}=\infty$, $i_{dc}^{\star}\leftarrow i_{dc}$
5:  repeat
6:    $\Omega_{\textrm{sca}}^{\star}\leftarrow\Omega_{\textrm{sca}}$, solve (31) to obtain $\boldsymbol{\Theta}$
7:    $l\leftarrow l+1$, $\mathbf{\Theta}^{(l)}\leftarrow\boldsymbol{\Theta}$, $\sigma\leftarrow\min\{1,1.5\sigma\}$
8:  until $\left\lVert 1-{\Omega_{\textrm{sca}}^{\star}}/{\Omega_{\textrm{sca}}}\right\rVert\leq\upsilon$
9:   $h_{n}=\mathbf{h}_{R,n}^{T}\mathbf{\Theta}^{(l)}{\mathbf{h}_{I,n}}$ and run Algorithm 1 to update $s_{n}^{(l)}$
10:   Compute $i_{dc}$ using (II-B)
11:until $\left\lVert 1-i_{dc}^{\star}/i_{dc}\right\rVert\leq\upsilon$
12:Run Algorithm 3 to obtain $\mathbf{\Theta}^{*}$
III-D Iterative Beamforming Optimization
@@
The proof is provided in Appendix B. By leveraging Proposition 2, one can linearize $h_{n}$ using the Neumann approximation and iteratively improve the approximation. Thus, by defining an auxiliary variable $\mathbf{\Omega}$ as the increment to the impedance matrix $\mathbf{Z}^{(l)}$ at the $l$th iteration of the Neumann approximation, we can write
@@
By leveraging the same matrix properties provided in Appendix A, we proceed by rewriting
$h_{n}$ $\approx{\mathbf{a}_{n}^{(l)}}^{T}(\mathbf{I}_{M}\ -\mathbf{\Omega}\mathbf{A}^{(l)})\mathbf{b}_{n}^{(l)}$
@@
Proposition 3 and Lemma 2 ensure that substituting (36) into (32a) leads to maximizing a convex objective function w.r.t. $\boldsymbol{\omega}$, which can be approximated similar to the waveform optimization procedure for fixed $h_{n}$. Specifically, the convex objective can be approximated using its first-order Taylor expansion at each local point and iteratively optimized. Thus, we proceed by writing the first-order Taylor’s coefficient as
@@
where (40c) ensures that $\mathbf{\Omega}$ has the same zero entries as $\mathbf{Z}$ depending on the structure of BD-RIS. Note that for a fully connected BD-RIS, constraint (40c) can be removed.
Let us proceed by defining $\bar{M}=M(M+1)/2$ and writing the Lagrangian of (40) as
It is obvious that by forcing $\boldsymbol{\omega}]_{r}=0,\forall r\in\mathcal{D}_{\mathcal{G}}$, we have $\bar{\lambda}_{r}^{(l)}[\boldsymbol{\omega}]_{r}=0,\forall r\in\mathcal{D}_{\mathcal{G}}$. Meanwhile, for $r\notin\mathcal{D}_{\mathcal{G}}$, the derivative of the Lagrangian function is given by
We know that the solution must satisfy KKT conditions [boyd2004convex]. Hence, by setting the derivative to zero, we can write the closed-form expression of the solution at the $l$th iteration as
$[\boldsymbol{\omega}^{*}]_{r}$ $=\begin{cases}0,&r\in\mathcal{D}_{\mathcal{G}},\\ {[\mathbf{u}^{(l)}]_{r}}/{2\nu_{r}^{(l)}},&\text{otherwise},\end{cases}.$ (43)
Then, using the complementary slackness for the dual variable $\nu_{r}^{(l)}$, and leveraging (43) and the fact that $\nu_{r}^{(l)}\geq 0,\forall r$, we have
$\nu_{r}^{(l)}$ $=\begin{cases}0,&r\in\mathcal{D}_{\mathcal{G}},\\ \left\lVert[\mathbf{u}^{(l)}]_{r}\right\rVert/2{\tau^{(l)}},&\text{otherwise}.\end{cases}$ (44)
Moreover, we introduce $\rho_{\omega}\leq 1$ to control the convergence of the solution, leading to writing the solution at the end of the $l$th iteration as
@@
which can be leveraged to solve (40) iteratively. Moreover, it was verified that $\mathrm{Vec}(\mathbf{\Omega}^{(l+1)})=\mathbf{P}\boldsymbol{\omega}^{(l+1)}$; thus, $\mathbf{\Omega}^{(l+1)}=\mathrm{Vec}^{-1}_{M,M}(\mathrm{Vec}(\mathbf{\Omega}^{(l+1)}))$ can be used for updating $\mathbf{Z}$ as
@@
Algorithm 5 illustrates the proposed iterative optimization method relying on alternating optimization, SCA, KKT conditions, and the Neumann approximation for BD-RIS-aided WPT (IT-BDRIS) given $\mathcal{D}_{\mathcal{G}}$. First, we initialize the variables such that the impedance matrix is set to the reference impedance value, while (15) is used to initialize $s_{n}$. Then, the alternating optimization is done through lines 6-16. Specifically, each alternating optimization iteration consists of iteratively updating $\mathbf{Z}$ given fixed $s_{n},\forall n$ in lines 8-13, followed by obtaining $s_{n},\forall n$ using Algorithm 1. This procedure is repeated iteratively until convergence.
The solution of the iterative beamforming design obtained in Algorithm 5 lines 8-13 converges to a stationary point of the approximate problem (40) (the proof is similar to Theorem 1). Although the solution does not necessarily represent a stationary point of the original problem (32), one can claim that the solution is close to a stationary point of (32) if the Neumann series approximation is tight. Algorithm 5 ensures the tightness of the Neumann series approximation by setting $\gamma\ll 1$ in 39, which leads to satisfying (34) [direnzomutual2, direnzoclerckxmutual, stewart1998matrix]. Following Theorem 1, Algorithm 5 provides an approximate feasible solution for (5).
III-E Extension to Scenarios with a Direct ET–ER Link
We now discuss the extension of the proposed algorithms to scenarios where a direct transmitter-receiver link is present. Let us denote the direct component as ${h}_{D,n}$, leading to the cascade channel being formulated as $h_{n}={h}_{D,n}+\mathbf{h}_{R,n}^{T}\mathbf{\Theta}{\mathbf{h}_{I,n}}$. This does not change the procedure of the waveform optimization algorithm since the cascade channel is an input to Algorithm 1. For SDR and SDP-based beamforming, we can rewrite (17) as $h_{n}=g{h}_{D,n}+\mathbf{a}_{n}^{T}\boldsymbol{\theta}$, where $g$ is an auxiliary variable. Then, $h_{n}$ can be reformulated as $h_{n}=\mathbf{\hat{a}}_{n}^{T}\boldsymbol{\hat{\theta}}$, where $\boldsymbol{\hat{\theta}}=[\boldsymbol{\theta},g]$ and $\mathbf{\hat{a}}_{n}=[\mathbf{a}_{n},{h}_{D,n}]$ are the concatenated vectors. Next, Algorithm 2 can be used to obtain $\boldsymbol{\theta}^{*}$, and the final solution can be computed by removing the last element in $\boldsymbol{\theta}^{*}/g$. Note that IT-BDRIS and SCA-BDRIS beamforming can be directly applied to the cases with a direct link only by substituting the new cascade channel (with a direct link).
@@ III-E Extension to Scenarios with a Direct ET–ER Link
III-F Complexity Analysis
Herein, we discuss the time complexity of the proposed algorithms in detail:
IT-WF: The time complexity at each optimization iteration is $\mathcal{O}(N)$, which corresponds to computing $\{s_{n}\}$.
SDR-BDRIS: The number of variables in (27) is ${\bar{M}(\bar{M}+1)}/2$, while the rest of the entries in $\mathbf{X}$ are determined according to the Hermitian structure, and the size of these Hermitian matrix sub-space is ${\bar{M}}^{2}$. Additionally, the number of constraints scales with $\bar{M}^{2}$ in (27). Interior-point methods require $\mathcal{O}(\sqrt{n})$ iterations, where $n$ is the number of constraints, and each iteration costs $\mathcal{O}\!\big(\max\{n\bar{M}^{3},\,n^{2}\bar{M}^{2},\,n^{3}\}\big)$ [boyd2004convex]. Thus, the per-iteration complexity is $\mathcal{O}(\bar{M}^{6})$, and the overall complexity grows polynomially with $\bar{M}$ in addition to the $\mathcal{O}(N)$ cost of IT-WF. Thus, SDR-BDRIS is significantly heavier than IT-BDRIS for large BD-RIS sizes.
IT-BDRIS: In the most time-consuming case, i.e., fully connected impedance network, [^5] Notice that the increased complexity of optimizing BD-RIS also leads to higher processing power consumption for computing the corresponding scattering matrix, which should be considered in practical implementations. $\boldsymbol{\omega}$ includes $\bar{M}$ non-zero entries, leading to $\mathcal{O}(\bar{M})$ complexity for obtaining the solution and the dual variables. Combining this with the negligible complexity for obtaining $\mathbf{Z}^{(l+1)}$ leads to $\mathcal{O}(\bar{M})$ time complexity for each optimization iteration. Thus, the proposed IT-BDRIS is much more efficient than the traditional solver-based SCA.
SCA-BDRIS: Each inner problem in (31) involves an $M\times M$ matrix variable and a fixed small number of linear matrix inequalities; thus, the per-iteration complexity is $\mathcal{O}(M^{3})$. Since several outer SCA iterations are required, the overall complexity is polynomial in $M$, higher than IT-BDRIS but substantially lower than the SDR-based formulation.
@@ IV Numerical Analysis
Figure 2: The simulation layout.
Figure 3: The convergence performance of (a) IT-WF (left), (b) IT-BDRIS (middle-left), (c) SCA-BDRIS (middle-right), (d) SDR-BDRIS (right). We adopt a random channel realization and assume a fully connected BD-RIS.
In this section, we evaluate the system’s performance in a WiFi-like scenario with $f_{c}=2.4$ GHz, a bandwidth of $\textrm{BW}=10$ MHz, and $\Delta f=\textrm{BW}/N$. We consider the path gain due to large-scale fading at a distance $d$ to be $L_{0}d^{-2}$, where $L_{0}=-40\penalty 10000\ \mathrm{dB}$ is the reference path power gain at a 1 m distance. The BD-RIS elements are positioned on a uniform planar array with $\lambda/2$ as the distance between two adjacent elements, where $\lambda$ is the wavelength. For the sake of simplicity, we consider a symmetric simulation setup where the incident and reflective path lengths are $d=2$ m, while $\theta_{e}=\phi_{e}=\pi/6$ are the elevation and azimuth angles. The simulation layout is illustrated in Fig. 2. The number of iterations for Algorithm 3 and Gaussian randomization is set to 50000. In SCA-BDRIS, initial value of $\sigma$ is set to $10^{-5}$ and $\iota=1$. Finally, the rectifier parameters are selected based on the circuit simulations in [clerckx2018beneficial, clreckxWFdesign] for the small-signal regime such that $K_{2}=0.17$ and $K_{4}=957.25$.
@@
where $\kappa$ is the Rician factor and the $\mathbf{h}_{n}^{\text{LoS}}$ is modeled using the far-field LoS channel for uniform planar arrays with elevation and azimuth angles as inputs. The NLoS part is modeled with Rayleigh fading considering $L=18$ delay taps with realizations following a circularly symmetric complex Gaussian distribution with a random power $p_{l}$, such that $\sum_{l=1}^{L}p_{l}=1$. Furthermore, we introduce an additional parameter $\alpha$ for the NLoS part such that the coherence bandwidth is $B_{c}=\alpha\textrm{BW}$. Hereby, the delay spread is $\varsigma=1/B_{c}$, while the maximum delay spread is considered to be $\varsigma_{m}=2\varsigma$. Then, the tap delays $t_{l},\forall l$ are generated uniformly spaced up to $\varsigma_{m}$, and the frequency response of each tap at the $n$th sub-carrier is multiplied by its corresponding delay term $e^{j2\pi f_{n}t_{l}}$. Hereby, selecting a small $\alpha$ such that $B_{c}$ is much smaller than BW leads to more frequency selectivity of the channel. Nevertheless, a large $\alpha$ leads to $B_{c}\gg\textrm{BW}$, which means all delay taps can be received within the coherence bandwidth, leading to a frequency-flat channel.
Note that the transmitter-RIS and RIS-receiver distances are chosen such that the far-field communications assumption holds. Specifically, the incident and reflective path lengths are larger than the Fraunhofer distance[MyEBDMA]. [^6] The Fraunhofer distance is formulated as $2D^{2}/\lambda$, where $D$ is the antenna diameter.The results are averaged over 200 random channel realizations unless otherwise stated. To ensure the tractability of the nonlinear rectenna model and maintain the validity of the fourth-order approximation, we consider up to eight subcarriers, which is sufficient to capture the key performance trends and allow a fair comparison between D-RIS and BD-RIS architectures.